# Advancing Question Answering on Handwritten Documents: A State-of-the-Art Recognition-Based Model for HW-SQuAD

### v1

@@ Abstract
Question-answering handwritten documents is a challenging task with numerous real-world applications. This paper proposes a novel recognition-based approach that improves upon the previous state-of-the-art on the HW-SQuAD and BenthamQA datasets. Our model incorporates transformer-based document retrieval and ensemble methods at the model level, achieving an Exact Match score of 82.02% and 92.55% in HW-SQuAD and BenthamQA datasets, respectively, surpassing the previous best recognition-based approach by 10.89% and 26%. We also enhance the document retrieval component, boosting the top-5 retrieval accuracy from 90% to 95.30%. Our results demonstrate the significance of our proposed approach in advancing question answering on handwritten documents. The code and trained models will be publicly available to facilitate future research in this critical area of natural language.
@@ 3 Method
In our proposed architecture, we improve the Document Retriever as well as the Document Reader parts. For document retrieval, we use additional preprocessing techniques, such as the Sentence transformer and TF-IDF. We used an ensemble of two large BERTs and one DeBERTa model for the Document Reader part. Our proposed architecture significantly improves document retrieval and reader performance.
@@ 3. Transformer Encoding:
In addition to the integration of sentence transformers, we employ advanced NLP preprocessing techniques to further enhance the accuracy of the document retrieval process. These techniques include tokenization and stemming. Tokenization involves breaking down the text into individual words or tokens. Stemming reduces words to their base or root form, helping to normalize the text and handle variations of the same word. By applying these preprocessing steps, we aim to focus on the most informative terms and improve the precision of the retrieved documents.
@@ 3.2 Document Reader Module
For the end-to-end recognition model, we propose an ensemble approach that combines the strengths of multiple extractive question answering models. Our ensemble consists of two BERT models with different initializations and one DeBERTa large model. BERT (Bidirectional Encoder Representations from Transformers) and DeBERTa (Decoding-enhanced BERT with Disentangled Attention) are both transformer-based models that have achieved state-of-the-art performance on various natural language processing tasks, including question answering.
Given an input sequence $x=[x_{1},x_{2},\dots,x_{n}]$, each model generates a sequence of hidden representations $H=[h_{1},h_{2},\dots,h_{n}]$ through multiple layers of self-attention and feed-forward networks:
$H_{\text{BERT}1}$ $=\text{BERT}1(x)$ (1)
$H_{\text{BERT}2}$ $=\text{BERT}2(x)$
$H_{\text{DeBERTa}}$ $=\text{DeBERTa}(x)$
For the question answering task, each model is fine-tuned to predict the start and end positions of the answer span within the context. The probability of a token $x_{i}$ being the start or end of the answer is computed using a softmax function over the hidden representations:
$P_{\text{start}}^{(m)}(i)=\frac{\exp(W_{\text{start}}^{(m)}\cdot h_{i}^{(m)})}{\sum_{j=1}^{n}\exp(W_{\text{start}}^{(m)}\cdot h_{j}^{(m)})}$ (2)
$P_{\text{end}}^{(m)}(i)=\frac{\exp(W_{\text{end}}^{(m)}\cdot h_{i}^{(m)})}{\sum_{j=1}^{n}\exp(W_{\text{end}}^{(m)}\cdot h_{j}^{(m)})}$
where $W_{\text{start}}^{(m)}$ and $W_{\text{end}}^{(m)}$ are learnable weight matrices for model $m\in{\text{BERT}_{1},\text{BERT}_{2},\text{DeBERTa}}$.
To combine the predictions from the three models, we employ an ensemble approach. The ensemble predictions are obtained by taking the union of the predicted answers from all models for each question. Let $A_{1},A_{2},A_{3}$ be the sets of predicted answers from the two BERT models and the DeBERTa model, respectively. The ensemble prediction $A_{\text{ensemble}}$ for a given question is computed as:
$A_{\text{ensemble}}=A_{1}\cup A_{2}\cup A_{3}$ (3)
The ensemble approach allows us to leverage the individual strengths of each model and potentially improve the overall performance compared to using a single model.
To evaluate the performance of the ensemble model, we use a custom evaluation function that calculates the number of correct, incorrect, and similar answers based on the ground truth. The function compares the predicted answers with the ground truth answers and categorizes them as follows:
Correct: If any of the predicted answers match the ground truth exactly. Similar: If any of the predicted answers have a significant overlap with the ground truth. Incorrect: If none of the predicted answers match or have significant overlap with the ground truth.
The evaluation function also keeps track of the corresponding question and answer texts for each category.
Figure 3: Comparison of correct, similar and incorrect among all the models.
@@ 4 Experimental setup
We evaluate the performance of our proposed model by utilizing the BenthamQA and HW-SQuAD datasets. By combining the OCR texts, we have created the context for our recognition-based pipeline. The question and the answer are derived from these two datasets. Following the extraction of context, questions, and answers, we implemented the basic preprocessing approach. Afterwards, we implemented our proposed Document Retriever. In order to implement the Recognition-based model, we converted the output to SQuAD dataset format. The HW-SQuAD dataset contains over 84,000+ pairs of QA pairs, while the Bentham-QA dataset contains 200 pairs. For training, validation, and testing, we have implemented the exact same split ratio as in [3]. We divided the Bentham-QA dataset into 80, 10, and 10 ratios for training, validation, and testing.
Our training was conducted using four 2080 ti Nvidia RTX graphics cards. AdamW is employed for BERT large, while Adam Optimizer is employed for the DeBERTa model. We employed two evaluation metrics for Document Reader: the F1 score and the EM [1][3]. The F1 score is the harmonic mean of precision and recall:
@@
Document Retriever was assessed based on the top 5-accuracy scores, as it is implemented in [3]. Top 5 accuracy is calculated as:
where:
  • $N$: Total number of instances
@@ 5.2 Impact of applying Ensemble method in Document Reader
Our proposed model demonstrates superior performance across both datasets. On the HW-SQuAD dataset, it achieves an F1 score of 90.10% and an Exact Match of 82.02%, representing a significant improvement over the baseline model’s 76.82% F1 and 70.73% EM. Similarly, on the BenthamQA dataset, our model attains a 93% F1 score and 90.42% EM, substantially outperforming the baseline’s 78.41% F1 and 66% EM.
@@ 5.3 Analysis of predicted results of our proposed model
In this section, we analyze the effect of applying the ensemble method to our proposed model. We have taken a few questions from the HW-SQuAD and BenthamQA datasets and compared the results in Fig. 5 and 6. respectively. In the HW-SQuAD dataset, the ensemble model consistently corrects predictions where the single model falters. For instance, in responses requiring specific information such as“divide to form new pyrenoids or be produced de novo,” the single model inaccurately predicts "some of their offspring." In contrast, the ensemble model accurately matches the ground truth. This precision is also evident in questions requiring detailed understanding, such as predicting “along the plant cells cell wall” instead of “under intense light.” Additionally, for broader categorical answers like “humid subtropical,” the ensemble model refines the single model’s vague prediction of "warm" to the precise ground truth. This meticulous correction and alignment with the expected answers underline the ensemble model’s enhanced interpretative capabilities. The effectiveness of the ensemble model is equally prominent in the BenthamQA dataset. The ensemble model not only rectifies broader, less accurate predictions of the single model but also ensures precision in complex queries. For example, it refines the single model’s generic prediction of “Offences against Property Theft” to the specific term “Embezzlement,” demonstrating its ability to grasp nuanced legal terminology. Moreover, for questions requiring multiple valid responses, such as listing various historical manufacturers or recognizing titles like “Lord Pelham,” the ensemble model accurately captures all relevant details, showcasing its comprehensive understanding and reliability.
@@ 5.4 Ablation Studies
To evaluate the effectiveness of different components in our proposed model, we conducted ablation studies on the HW-SQuAD dataset, as shown in Table 3.
First, we examined the impact of our document retrieval enhancements. The baseline model [3] achieved an F1 score of 76.82 and an Exact Match of 70.2. By adding the Sentenece Transformer (ST) module to the retrieval pipeline, our TF-IDF + ST + BERT model improved the F1 score to 83.20 and the Exact Match to 71.33. This demonstrates the effectiveness of incorporating semantic information for retrieving more relevant documents.
Next, we analyzed the performance of our complete proposed pipeline, which includes an ensemble of multiple reader models in addition to the retrieval enhancements. The TF-IDF + ST + Ensemble model achieved an impressive F1 score of 90.10 and an Exact Match of 82.02, significantly outperforming the baseline. These results highlight the importance of both the improved document retrieval and the ensemble strategy in our proposed approach.
@@ 6 Conclusion
Table 3: Ablation studies of Our Model.
Model Details HW-SQuAD
F1 Exact Match
a. TF-IDF + BERT [3] 76.82 70.2
b. TF-IDF + ST + BERT (proposed) 83.20 71.33
c. TF-IDF + ST + Ensemble (proposed) 90.10 82.02

### v2

@@ Abstract
Question-answering on handwritten documents is a challenging task with numerous real-world applications. This paper proposes a novel recognition-based approach that improves upon the previous state-of-the-art on the HW-SQuAD and BenthamQA datasets. Our model incorporates transformer-based document retrieval and ensemble methods at the model level, achieving an Exact Match score of 82.02% and 69% in HW-SQuAD and BenthamQA datasets, respectively, surpassing the previous best recognition-based approach by 10.89% and 3%. We also enhance the document retrieval component, boosting the top-5 retrieval accuracy from 90% to 95.30%. Our results demonstrate the significance of our proposed approach in advancing question answering on handwritten documents. The code and trained models will be publicly available to facilitate future research in this critical area of natural language.
@@ 3 Method
In our proposed architecture, we improve the Document Retriever as well as the Document Reader parts. For document retrieval, we use additional preprocessing techniques, such as the Sentence transformer and TF-IDF. We used an ensemble of two large BERTs and one DeBERTa model for the Document Reader part. Our proposed architecture significantly improves document retrieval and reader performance. Figure 3 depicts the entire architecture of our proposed model.
@@ 3. Transformer Encoding:
In addition to the integration of sentence transformers, we employ NLP preprocessing techniques to further enhance the accuracy of the document retrieval process. These techniques include tokenization and stemming. Tokenization involves breaking down the text into individual words or tokens. Stemming reduces words to their base or root form, helping to normalize the text and handle variations of the same word. By applying these preprocessing steps, we aim to focus on the most informative terms and improve the precision of the retrieved documents.
Algorithm 2 Ensemble Approach for Document Reader Module
1:Initialize three models: Two BERT models (BERT1, BERT2) and one DeBERTa model
[…]
$\mathbf{P_{\text{start}}^{(m)}(i)=\text{softmax}(W_{\text{start}}^{(m)}\cdot h_{i}^{(m)})}$
$\mathbf{P_{\text{end}}^{(m)}(i)=\text{softmax}(W_{\text{end}}^{(m)}\cdot h_{i}^{(m)})}$
•  where $W_{\text{start}}^{(m)}$ and $W_{\text{end}}^{(m)}$ are learnable weight matrices •  Update model parameters to minimize the loss function end for
3:Testing and Evaluation Phase: for each question in the dev and test set do •  Generate predictions using each fine-tuned model:
[…]
$\mathbf{A_{3}=\text{DeBERTa}(\text{question},\text{context})}$
•  Compute ensemble prediction:
$\mathbf{A_{\text{ensemble}}=A_{1}\cup A_{2}\cup A_{3}}$
•  Evaluate ensemble performance: •  Compare $A_{\text{ensemble}}$ with ground truth answer •  Categorize prediction as: •  a. Correct: Exact match with ground truth •  b. Similar: Significant overlap with ground truth •  c. Incorrect: No match or significant overlap end for
4:Compute overall performance metrics (e.g., accuracy, F1 score)
@@ 3.2 Document Reader Module
We propose an ensemble approach for the Document Reader module (depicted in Figure 2) that combines the strengths of multiple extractive question-answering models. Our ensemble consists of two BERT models with different initializations and one large DeBERTa model. BERT (Bidirectional Encoder Representations from Transformers) and DeBERTa (Decoding-enhanced BERT with Disentangled Attention) are transformer-based models that perform state-of-the-art natural language processing tasks, including question answering. The Algorithm 2 describes the workings of the BERT and DeBERTa models and then describe how we ensemble them for evaluating dev and test sets.
This algorithm presents the steps for an ensemble approach for the Document Reader module in a question-answering system. It combines two large BERT (Bidirectional Encoder Representations from Transformers) models and one large DeBERTa (Decoding-enhanced BERT with Disentangled Attention) model to leverage their strengths and potentially improve overall performance. The algorithm is divided into two main phases:
  1. 1.
    Training Phase: Each model (BERT1, BERT2, and DeBERTa) is fine-tuned separately on the question-answering task. It involves generating hidden representations, computing start and end probabilities for answer spans, and updating model parameters to minimize the loss function.
  2. 2.
    Testing and Evaluation Phase: During inference, for each question in the test and validation set, predictions are generated using all three fine-tuned models. The ensemble prediction is formed by taking the union of predicted answers from all models. The algorithm then evaluates the ensemble performance by comparing predictions with ground truth answers and categorizing them as correct, similar, or incorrect.
This ensemble method aims to enhance the accuracy and reliability of the question-answering system by combining multiple state-of-the-art models. The diversity in predictions can lead to more robust results, potentially capturing nuances that individual models might miss.
@@ 4 Experimental setup
We evaluate the performance of our proposed model by utilizing the BenthamQA and HW-SQuAD datasets. By combining the OCR texts, we have created the context for our recognition-based pipeline. The question and the answer are derived from these two datasets. Following the extraction of context, questions, and answers, we implemented the basic preprocessing approach. Afterwards, we implemented our proposed Document Retriever. In order to implement the Recognition-based model, we converted the output to SQuAD dataset format. The HW-SQuAD dataset contains over 84,000+ pairs of QA pairs, while the Bentham-QA dataset contains 200 pairs. For training, validation, and testing, we have implemented the exact same split ratio as in [3]. WE have fintuned our model on HW-SQuAD dataset. The BenthamQA dataset remains unchanged and only used for evaluation.
Our training was conducted using four 2080 ti Nvidia RTX graphics cards. AdamW is employed for BERT large, while Adam Optimizer is employed for the DeBERTa model.
4.1 Evaluation Metrics
We utilized the top 5-accuracy scores for evaluating the Document Retriever Module, as it is implemented in [3]. Top 5 accuracy is calculated as:
where:
  • $N$: Total number of instances
We employed two evaluation metrics for Document Reader: the F1 score and the EM [1][3]. The F1 score is the harmonic mean of precision and recall:
@@ 5.2 Impact of applying Ensemble method in Document Reader
Our proposed model demonstrates superior performance across both datasets. On the HW-SQuAD dataset, it achieves an F1 score of 90.10% and an Exact Match of 82.02%, representing a significant improvement over the baseline model’s 76.82% F1 and 70.73% EM. Similarly, on the BenthamQA dataset, our model attains a 81.57% F1 score and 69% EM, substantially outperforming the baseline’s 78.41% F1 and 66% EM.
@@ 5.3 Analysis of predicted results of our proposed model
In this section, we analyze the effect of applying the ensemble method to our proposed model. We have taken a few questions from the HW-SQuAD and BenthamQA datasets and compared the results in Fig. 5 and 6. respectively. In the HW-SQuAD dataset, the ensemble model consistently corrects predictions where the single model falters. For instance, in responses requiring specific information such as “divide to form new pyrenoids or be produced de novo,” the single model inaccurately predicts “some of their offspring.” In contrast, the ensemble model accurately matches the ground truth. This precision is also evident in questions requiring detailed understanding, such as predicting “along the plant cells cell wall” instead of “under intense light.” Additionally, for broader categorical answers like “humid subtropical,” the ensemble model refines the single model’s vague prediction of “warm” to the precise ground truth. This meticulous correction and alignment with the expected answers underline the ensemble model’s enhanced interpretative capabilities. The effectiveness of the ensemble model is equally prominent in the BenthamQA dataset. The ensemble model not only rectifies broader, less accurate predictions of the single model but also ensures precision in complex queries. For example, it refines the single model’s generic prediction of “Offences against Property Theft” to the specific term “Embezzlement,” demonstrating its ability to grasp nuanced legal terminology. Moreover, for questions requiring multiple valid responses, such as listing various historical manufacturers or recognizing titles like “Lord Pelham,” the ensemble model accurately captures all relevant details, showcasing its comprehensive understanding and reliability.
@@
Figure 6: Comparison of correct, similar and incorrect among all the models. The addition of our proposed methodology improves the no of corrects and reducing the no of similar and incorrect result. We have depicted the how addition of each technique over the old model improves the performance
Table 3: Ablation studies of Our Document Retriever Module.
Model Details Top-5 accuracies
HW-SQuAD BenthamQA
a. TF-IDF 90.2 98.5
b. TF-IDF + pre-processing 90.8 98.94
c. TF-IDF + pre-processing + ST 95.30 100
Table 4: Ablation studies of Our Document Reader Module on HW-SQuAD
Model Details HW-SQuAD
F1 Exact Match
a. TF-IDF + BERT [3] 76.82 70.2
b. TF-IDF + BERT (Our implementation) 77.46 63.21
c. TF-IDF + preprocessing + BERT (proposed) 81.18 68.35
d. TF-IDF + preprocessing + ST + BERT (proposed) 83.20 71.33
e. TF-IDF + preprocessing + ST + Ensemble (proposed) 90.10 82.02
Table 5: Ablation studies of Our Document Reader Module on BenthamQA
Model Details BenthamQA
F1 Exact Match
a. TF-IDF + BERT [3] 78.41 66
b. TF-IDF + BERT (Our implementation no finetuning) 64.75 53
c. TF-IDF + preprocessing + BERT (proposed) 73.96 59.2
d. TF-IDF + ST + preprocessing + BERT (proposed) 73.96 59.2
e. TF-IDF + ST + preprocessing + ensemble (proposed) 81.57 69.1
@@ 5.4 Ablation Studies
Tables 3 and 4 illustrate the ablation studies that were conducted on the HW-SQuAD and BenthamQA datasets to assess the efficacy of various components in our proposed model.
As an initial step, we investigated the impact of preprocessing and the Sentence Transformer on our Document Retriever Module. Using only the TF-IDF and basic preprocessing, the baseline model [3] achieved 90.2 top-5 accuracy. We have implemented advanced preprocessing techniques, resulting in a top-5 accuracy score of 90.8. Furthermore, we implemented the Sentence Transformer (ST), which has a significant effect on performance, resulting in a 5% increase. Additionally, in the BenthamQA dataset, improving the preprocessing results in an approximate 0.5% increase in top-5 accuracy, and when the ST is added to top of that, it reaches 100%.
Next we have done ablation studies for the Document Reader part for both dataset HW-SQuAD and BenthamQA. Fird we describe the effect on the HW-SQuAD dataset which is represented in Table 4. Here we have listed the result reported in [3] and the we have given our results. At first we have applied the single BERT model and we reached 63.21 Exact Match and 77.46 F1 score without any type of preprocessing. when we add our proposed preprocessing to the Retrieval pipeline the it reached 68.21 EM and 77.46 F1 score. By adding the Sentence Transformer (ST) module to the retrieval pipeline, the model gets improved and achieved F1 score of 83.20 and the Exact Match of 71.33. This demonstrates the effectiveness of incorporating semantic information for retrieving more relevant documents.
The ablation studies in the BenthamQA dataset gets presented in the Table 5. Like the Table 4 we first give the result in the HW-SQuAD paper. Next we have given the result of our implementation. Then we have add our proposed techniques and how it improves the performance. When we slightly finetuned and add the preprocessing the model scored 82.97 EM. When we add the ST to it the EM improves further to 86.17%. Like in HW-SQuAD Semantic information in time of retrieval improves the performance of Document reader significantly.
Next, we analyzed the performance of our ensemble approach, which includes three extractive QA models (two BERT and one DeBERTa large) in addition to the retrieval enhancements. The TF-IDF + ST + Ensemble model achieved an impressive F1 score of 90.10 and an Exact Match of 82.02, significantly outperforming the baseline. Also in BenthamQA dataset, when we add the ensemble method we reach over 90% EM and 96.12% F1 score. These results highlight the importance of both the improved document retrieval and the ensemble strategy in our proposed approach.